OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 4 and 9. What is the length of the altitude?

OpenStudy (anonymous):



OpenStudy (anonymous):

m^2=9*4=36 m=6

OpenStudy (anonymous):

Oh, thankss(: